OpenStudy (anonymous):

Plan A cost $10 plus $0.13 per minute. Plan B cost $15 plus $0.09 per minute. What amount of calling do the two plans cost the same ? And what is the cost when the two plans cost the same ?

OpenStudy (anonymous):

Let x be the number of minutes when both of the plan costs are equal. Cost is in pennies. Solve the following for x: 1000 +13x = 1500 + 9x x = 125 minutes Replace x with 125 minutes. {10+0.13x,15+0.09x} /. x->125 minutes {$26.25, $26.25}

OpenStudy (anonymous):

Omg thank youu ! I've been trying to figure this out for like a hour , I wasn't thinking of adding the $10 and $15 I was only adding the $0.13 and $0.09 for the minutes ! Haha thanks again .

OpenStudy (anonymous):

Your welcome.